5. (3 points) A mass-spring system is in SHM in the horizontal direction. If the mass of 1.2 kg, the spring constant is 15 N/m and the amplitude is 20.0 cm, what is the speed of the mass at the moment when the displacement is 10.0 cm. Hint: Speed is always expressed as a nonnegative number (positive number or zero)
Answer #1


angular frequency w = sqrt(k/m) = sqrt(15/1.2) = 3.54 /s

speed v = w*sqrt(A^2-x^2)

speed v = 3.54*sqrt(0.2^2-0.1^2) = 0.613 m/s <<<-----------ANSWER